OpenStudy (anonymous):

divide and check. HELP! x^4+x^3/x

OpenStudy (anonymous):

\[\frac{ x^{4}+x ^{3}}{ x}\]

geerky42 (geerky42):

Try to factor numerator. What do both terms in numerator have in common?

OpenStudy (anonymous):

x

geerky42 (geerky42):

Yup so factor it out.

OpenStudy (anonymous):

x^7

OpenStudy (anonymous):

x^7/x?

geerky42 (geerky42):

no, just factor it out, like for example: factor a out from a² + a³ \(\Rightarrow\) a(a + a²)

geerky42 (geerky42):

Confused?

OpenStudy (anonymous):

x(x^3+x^2)?

OpenStudy (anonymous):

beyond!! math is not my strongest suit. :(

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@bahrom7893 would you be able to help me? I'm stuck

OpenStudy (jim766):

\[\frac{ x^3(x+1) }{ x } \]

OpenStudy (anonymous):

@Jim766 how so tho?

OpenStudy (anonymous):

(x^4+x^3)/x = x^3 + x^2 how did I end up with this? agh! I'm confused! :(

OpenStudy (jim766):

x^4 = x*x*x*x + x^3 = x*x*x they both have x^3 in common

OpenStudy (jim766):

then it reduces to x^2(x+1)

OpenStudy (anonymous):

so my answer would be x^2(x+1)/x?

OpenStudy (jim766):

\[x^4 + x^3\] greatest common factor is \[x^3\] now divide x^4+ x^3 by x^3 to factor it out x^4 + x^3 --- ---- = x^1 + 1 x^3 x^3 x^3(x + 1) = x^4 + x^3